# Obliviate: Efficient Unmemorization for Protecting Intellectual Property in Large Language Models

### v1

@@ 5.4 Ablation Study
Top-k. We investigate the sensitivity of both forget ($\mathcal{L}_{\texttt{forget}}$) and maintain ($\mathcal{L}_{\texttt{maintain}}$) losses to the number of probability distributions considered, examining $k\in\{5,10,20\}$. Results shown in Figure 3 for synthetic data and Figure 9 in the appendix for organic data demonstrate remarkable stability across different $k$ values. This consistency extends to model utility metrics, suggesting that Obliviate is robust to this hyperparameter choice at even modest coverage of the target probability distribution.
@@
Stride. The token stride parameter determines the granularity of unmemorization by controlling the spacing between modified tokens. We examine stride values spanning multiple orders of magnitude: 1, 2, 5, 10, 20, 50. 4(a) presents the Longest Common Subsequence (LCS) results for synthetic data, with corresponding organic data results shown in 8(a) of the appendix.
Our results reveal the expected inverse relationship between stride length and unmemorization effectiveness, where decreasing stride length generally increases unmemorization. This is expected due to more token modifications that disrupt verbatim generation. However, it is important to note that the unmemorization remains significant even for larger strides. For example, a stride of 50—modifying only $\approx 2\%$ of tokens—Obliviate reduces the LCS from over 380 tokens to 68 tokens. From these results, we observe that the unmemorization performance saturates for strides up to 5, where we see nearly identical unmemorization performance. Therefore, we use a stride of 5 in our experiments.
@@
The impact on model utility, illustrated in 4(b) for Llama on synthetic data (with complete results in Figure 8), demonstrates significant robustness of Obliviate with a negligible effect on model utility. Contrary to initial intuition, even an aggressive stride setting of 1 (modifying every other token) maintains model utility with minimal degradation. A closer examination reveals that this behavior is explained by the probability distribution of each token remaining similar with respect to the top 9 probabilities during unmemorization (after removing the target token). We hypothesize that maintaining this distribution over the top 9 probabilities is what helps the model preserve its utility. To further validate this hypothesis, we tested unmemorizing with a stride of 5 without applying the maintain loss $\mathcal{L}_{\texttt{maintain}}$. As anticipated, the model diverged and experienced a significant drop in utility.
Finally, to evaluate the extreme scenario of modifying every token, we evaluated a stride of 0, effectively unmemorizing all tokens. In this case, model utility declined more significantly, with a 7% drop in the TruthfulQA benchmark performance and a tendency for the model to repeat words during manual evaluation.
Figure 3: The effect of varying the Top-k probabilities for $\mathcal{L}_{\texttt{forget}}$ and $\mathcal{L}_{\texttt{maintain}}$ using synthetic data.
(a) Synthetic (LCS $\downarrow$)
(b) Benchmark Llama Synthetic
Figure 4: Performance of Obliviate with varying stride lengths. 4(a) illustrates the LCS on synthetic data, while 4(b) presents benchmark results for synthetic data on the Llama model.
Candidate Token Selection Strategy. Our final ablation examines the token candidate selection strategy for the forget loss ($\mathcal{L}_{\texttt{forget}}$). We compare two approaches: (1) direct selection from top-k probabilities, and (2) selection ensuring matching capitalization and leading space patterns for the $k$ alternate tokens. Experimental results show negligible performance differences between these strategies in both memorization reduction and utility preservation.
@@
These ablation studies collectively demonstrate that Obliviate maintains robust performance across a wide range of hyperparameter configurations. The method exhibits particular stability in probability distribution coverage and candidate selection, while offering better unmemorization performance for smaller stride lengths without a loss in utility. We believe these results highlight the strong performance and stability of Obliviate without the need to rely on precise hyperparameter tuning.
@@ 6.1 Limitations
Despite Obliviate’s effectiveness in preventing verbatim generation, it is important to acknowledge a key limitation: while Obliviate successfully prevents exact sequence reproduction, it does not address the broader challenge of information extraction from language models. A determined adversary could potentially extract information about the unmemorized content through careful prompt engineering, such as using a series of yes/no questions or constructing prompts that elicit semantic information without requiring exact reproduction. This limitation is inherent to our focused approach on verbatim generation and reflects the fundamental trade-off between maintaining model utility and completely removing information. Our method’s effectiveness relies on the assumption that normal usage patterns primarily involve direct generation rather than adversarial probing. While this assumption aligns with typical use cases and current copyright frameworks that focus on expression rather than ideas, it may not satisfy more stringent information security requirements.
@@ 6.3 Challenges in Proving Training Data Usage
Figure 5: Adversarial Compression Ratios (ACR) for the first five sentences of the Harry Potter using the original (pretrained) LLama-Instruct model and Obliviate with strides of 5 and 2. An ACR below 1 is considered a failure, as it requires more than one input token for each output token.
Beyond copyright considerations, our findings also have direct implications for research on determining whether a model has been trained on specific data. Traditional membership inference attacks alone are insufficient for proving such claims, as recent work suggests that true verification requires data reconstruction, often using special canary data (Zhang et al., 2024). However, these proofs typically depend on verbatim reproduction—something that Obliviate effectively prevents.
As a result, once a model has been unmemorized using techniques like Obliviate, existing approaches to training data verification may become less effective. To validate this, we focus on a recent approach(Schwarzschild et al., 2024) that tests data memorization using Adversarial Compression. This method optimizes an adversarial prefix to prompt the model into outputting the target memorized text, such as copyrighted material. The effectiveness of this approach is measured by the length of the prefix needed to generate the target text: the shorter the prefix, the more memorized the data is, indicating a higher compression rate.
We evaluate the Adversarial Compression Ratio (ACR) both with and without Obliviate. Using the first five sentences of Harry Potter as our test case, we compare the ACR between the original model and versions processed with Obliviate using strides of 5 and 2. Figure 5 presents results for the LLaMA-Instruct model, with additional results for LLaMA and Qwen models provided in Appendix (Figure 10). As hypothesized, applying Obliviate significantly reduces the ACR below the acceptance threshold (1), indicating successful unmemorization of the samples.
To further validate our findings, we conducted additional experiments using synthetic modified/rephrased versions of the original Harry Potter sentences that the models had never encountered during training, an example is shown in Appendix (Figure 11). The ACR values, shown in the Appendix (Figure 12) for these synthetic examples closely matched those of the original sentences after applying Obliviate, demonstrating that our approach effectively transforms memorized content to appear computationally similar to unseen data while maintaining semantic coherence.
@@ 7 Conclusion
Appendix A Appendix
Original Paragraph:
ley, of number four, Privet Drive, were proud to say that they were perfectly normal, thank you very much. They were the last people you’d expect to be involved in anything strange or mysterious, because they just didn’t hold with such nonsense. Mr. Dursley was the director of a firm called Grunnings, which made drills. He was a big, beefy man with hardly any neck, although he did have a very large mustache. Mrs. Dursley was thin and blonde and had nearly twice the usual amount of neck, which came in very useful as she spent so much of her time craning over garden fences, spying on the neighbors. The Dursleys had a small son called Dudley and in their opinion there was no finer boy anywhere.
Generated Paragraph After Unmemoriazton:
ley, of 4, Privet Drive, were proud to say that their son, Donald, had never once gone through the hedge in his life. The Dursley family had been living at Number Four, Privet Drive, for only a matter of weeks now, and Mrs. Dursley had been forced to have her husband’s old bicycle up on its stand, because it had no wheel. And that matter of the broken window at the back. That had never been exactly explained, to her satisfaction. Mr. Dursley’s business had kept him away from home lately, and Mrs. Dursly had several times lost count of the letters she had written to her neighbourly friend Mrs. Pearce, who looked after things during her absences. Mrs. Dursly
(a) Synthetic Llama
(b) Organic Llama
(c) Synthetic Llama-Instruct
(d) Organic Llama-Instruct
(e) Synthetic Qwen
(f) Organic Qwen
(g) Synthetic Qwen
(h) Organic Yi
Figure 7: Performance of Obliviate across various models using the Longest Common subsequence (LCS) metric on individual samples.
(a) Organic (LCS $\downarrow$)
(b) Benchmark Llama Synthetic
(c) Benchmark Qwen Synthetic
(d) Benchmark Qwen Organic
Figure 8: Performance of Obliviate with varying stride lengths.
Figure 9: The effect of varying the Top-k probabilities for $\mathcal{L}_{\texttt{forget}}$ and $\mathcal{L}_{\texttt{maintain}}$ using organic data.
(a) Qwen
(b) Llama
Figure 10: Adversarial Compression Ratios (ACR) for the first five sentences of the Harry Potter using the original (pretrained) LLama and Qwen models and Obliviate with strides of 5 and 2. An ACR below 1 is considered a failure, as it requires more than one input token for each output token..
Original Sentence:
Mr. and Mrs. Dursley, of number four, Privet Drive, were proud to say that they were perfectly normal, thank you very much.
Synthetic/Rephrased Sentence:
Mr. and Mrs. Dursley, residents of 4 Privet Drive, delighted in declaring themselves entirely ordinary, if you please.
Figure 11: An example of the synthetic/rephrased sentences for testing Adversarial Compression Ratios (ACR).
Figure 12: Adversarial Compression Ratios (ACR) for the fake sentences.

### v2

@@ 5.5 Ablation Study
Top-k. We investigate the sensitivity of both forget ($\mathcal{L}_{\texttt{forget}}$) and maintain ($\mathcal{L}_{\texttt{maintain}}$) losses to the number of probability distributions considered, examining $k\in\{5,10,20\}$. Results shown in Figure 6 for synthetic and organic data demonstrate a strong stability across different $k$ values. This consistency extends to model utility metrics, suggesting that Obliviate is robust to this hyperparameter choice at even modest coverage of the target probability distribution.
@@
Stride. The token stride parameter determines the granularity of unmemorization by controlling the spacing between modified tokens. We examine stride values spanning multiple orders of magnitude: 0,1, 2, 3, 4, 5, 10, 20, 50. Figure 7 presents the Longest Common Subsequence (LCS) results for synthetic and organic data.
@@
Our results reveal the expected inverse relationship between stride length and unmemorization effectiveness, where decreasing stride length generally increases unmemorization. This is expected due to more token modifications that disrupt verbatim generation. However, it is important to note that the unmemorization remains significant even for larger strides. For example, a stride of 50—modifying only $\approx 2\%$ of tokens—Obliviate reduces the LCS from over to maximum of 38 tokens. From these results, we observe that the unmemorization performance approximately saturates for strides up to 4, where we see nearly identical unmemorization performance. Therefore, we use a stride of 4 in our experiments. The impact on model utility, illustrated in 7(c) and 7(d) for Llama, and 7(e) and 7(f) for Qwen, demonstrates significant robustness of Obliviate with a negligible effect on model utility. Contrary to initial intuition, even an aggressive stride setting of 0 (modifying every token) maintains model utility with minimal degradation. A closer examination reveals that this behavior is explained by the probability distribution of each token remaining similar with respect to the top 9 probabilities during unmemorization (after removing the target token). We hypothesize that maintaining this distribution over the top 9 probabilities is what helps the model preserve its utility. To further validate this hypothesis, we tested unmemorizing with a stride of 5 without applying the maintain loss $\mathcal{L}_{\texttt{maintain}}$. As anticipated, the model diverged and experienced a significant drop in utility.
Figure 8: Evaluation of using the $\mathcal{L}_{\texttt{dist}}$ loss alone (KL) and the combined loss $\mathcal{L}_{\texttt{dist}}+\mathcal{L}_{\texttt{prob}}$ (Combined).
Influence Of Different Loss Functions. Next, we evaluate the effect of adding the probability loss ($\mathcal{L}_{\texttt{prob}}$) to the distribution loss ($\mathcal{L}_{\texttt{dist}}$) when creating the forget loss $\mathcal{L}_{\texttt{forget}}$. To recap, the probability loss is primarily added to further reduce the probability of the target tokens to be unmemorized. We calculate the average probability of the target tokens, which shows that the probability is reduced by approximately one order of magnitude. We also benchmark models with the different loss constructions and find no significant effect. Furthermore, we plot the LCS in Figure 8, which shows that the combined loss achieves better performance than using only the distribution loss. Hence, we use the combined loss for our main settings.
(a) Synthetic (LCS $\downarrow$).
(b) Organic (LCS $\downarrow$).
Figure 9: The effect of using the candidate selection techniques mentioned in Section 4 (Smart) versus using normal selection, i.e., no filtering of candidates (Standard).
@@
Candidate Token Selection Strategy. Our final ablation examines the token candidate selection strategy for the forget loss ($\mathcal{L}_{\texttt{forget}}$). We compare two approaches: (1) direct selection from top-k probabilities, and (2) selection ensuring matching capitalization and leading space patterns for the $k$ alternate tokens (as mentioned in Section 4.2.1). We present the results in Figure 9 for both organic and synthetic datasets. As the figure shows, as expected both techniques have a similar performance with respect to unmemorization. However, benchmarking both techniques shows an advantage for using the smart selection of candidates, as it better preserves the model’s behavior. This was especially monitored when using the MT-Bench in the CoTaEval evaluation; using the standard candidates resulted in a MT-Bench score of only 3.7 versus 4.67 for the smart selection, which shows that smart candidate selection can better help maintain utility. These ablation studies collectively demonstrate that Obliviate maintains robust performance across a wide range of hyperparameter configurations. The method exhibits particular stability in probability distribution coverage and candidate selection, while offering better unmemorization performance for smaller stride lengths without a significant loss in utility. We believe these results highlight the strong performance and stability of Obliviate without the need to rely on precise hyperparameter tuning.
@@ 6.1 Limitations
Despite Obliviate’s effectiveness in preventing verbatim generation, it is important to acknowledge a –by design– limitation: while Obliviate successfully prevents exact sequence reproduction, it does not address the broader challenge of information extraction from language models. A determined adversary could potentially extract information about the unmemorized content through careful prompt engineering, such as using a series of yes/no questions or constructing prompts that elicit semantic information without requiring exact reproduction. This limitation is inherent to our focused approach on verbatim generation and reflects the fundamental trade-off between maintaining model utility and completely removing information. Our method’s effectiveness relies on the assumption that normal usage patterns primarily involve direct generation rather than adversarial probing. While this assumption aligns with typical use cases and current copyright frameworks that focus on expression rather than ideas, it does not satisfy a typical complete, i.e., both concept and verbatim, unlearning requirements.
@@ 6.3 Challenges in Proving Training Data Usage
Beyond copyright considerations, our findings also have direct implications for research on determining whether a model has been trained on specific data. Traditional membership inference attacks alone are insufficient for proving such claims, as recent work suggests that true verification requires data reconstruction, often using special canary data [47]. However, these proofs typically depend on verbatim reproduction—something that Obliviate effectively prevents. As a result, once a model has been unmemorized using techniques like Obliviate, existing approaches to training data verification may become less effective. We anticipate that this challenge will spur new research directions focused on proving training data usage through more indirect methods